Perform the following structural analysis on the 60-ft. pedestrian walkway deck truss and floor system. Assume only service loads (no load factors or impact).

(Stringer Calculations – Pedestrian Loading)

1.      Calculate all dead and live loads.

2.      Calculate the maximum reactions for the exterior and interior stringers.

3.      Develop the shear and moment diagrams for the exterior and interior stringers.

-o ?? 12 (w) o (A CONCRETE CaRB wizxs T? 12-0 Trensses 2-0 AL MEmBes Teuss MemBERS ARE MADE 1 ASTm AnD9 GRaDE SD STEEL. E: 29000 K3
Answer #1

Dead ood Kailing/ 121 p/fReduryj om f.oad = 0.075 KSf load dnbnynhonv- 4 P2 C -12 R, R2 455 知んn hal C38Member AB BA BC св CD DC DE ED Distribution Factor Fixed End Moment 0.428571 0.571429 0.5 0.5 0.571429 0.428571 18.75 18.7518nd Shaon and bending for unit Raading Shringen Floon beame I S 1ら S IS IX 18. 1 2 烹 24-1 I 6、04 I 8 R. ig 2+ 24.10? 9107 t 2Chhennaリ > = 24 며 x0.455 16-07x0.638= 10-25 Kith [tntennal] S.F 9656 eMomal 5 819X 0.638 -3.76 Kir 5.893x6 49sChhennaリ > = 24 며 x0.455 16-07x0.638= 10-25 Kith [tntennal] S.F 9656 eMomal 5 819X 0.638 -3.76 Kir 5.893x6 49s